# thanks: present address

### v3

@@ I Introduction
Since the phenomenon of QMPE is a comparatively new area of research, it is important to explore the plausible connections of QMPE to other interesting features of quantum systems. In the present work, we would like to address some such connections. In particular, the study of exceptional points (EPs) and surfaces in non-Hermitian quantum systems and Liouvillian dynamics has caught a lot of attention in recent days Heiss12 ; Hashimoto15 ; Doppler16 ; Zhong19 ; McDonald20 ; Soleymani22 ; Ding22 ; Minganti19 ; Arkhipov20 ; Arkhipov21 ; Muller08 ; Naghiloo19 ; Saha23 ; Roccati23 along with their applications in quantum optics Miri19 ; Ozdemir19 . The occurrence of EPs is directly related to eigen-spectrum of a system and hence to its relaxation process. Therefore a natural curiosity would be to understand the role of EPs on the possibility of QMPE. In fact, the occurrence of double MPE (i.e. two intersections between trajectories) at an EP has already been pointed out for MPE in classical inertial suspensions Takada21a . Interestingly, the acceleration of relaxation process in open quantum systems by maximizing the gap between the zero mode and the slowest eigenmode at an EP, has recently been studied Zhou23 . In general, such speed up is desired for several application purposes and therefore it has been a subject of immense interest in both classical and quantum systems Lapolla20 ; Vu21 ; Kochsiek22 ; Ghimenti22 ; Bao22 ; Zhou23 . Since QMPE allows faster relaxation of initially higher distant trajectories, it is important to find criteria for QMPE and to understand how to utilize QMPE to achieve faster approach to steady state. Another intriguing feature in non-Hermitian quantum mechanical systems is the appearance of complex eigenvalues Bender2007 ; Moiseyev2011 ; Nakano2011 ; Hatano2015 ; Li2023 . It is noteworthy that the analysis of QMPE in the existing examples Nava19 ; Carollo21 ; Manikandan21 ; Ivander23 ; Chatterjee_2023 is based on purely exponential relaxation processes, neither the EP nor the oscillations have been considered yet. There is only an exception in Ref. Ares23 where more than one intersections between trajectories (two intersections to be precise, termed as revivals) has been observed in the entanglement entropy in the XXZ spin chain, although not discussed much and simply explained as a finite-size effect.
@@
The paper is arranged as follows. In Sec. II, we define the details of the model and the protocol used to study QMPE. The Sec. III describes different regions in the control parameter plane depending on the types of eigenvalues of the Lindbladian and the multiple QMPE is defined. In articular, we focus on regions with exceptional points and complex eigenvalues to discuss their effects on QMPE in the density matrix and observables such as energy, entropy, Kullback-Leibler (KL) divergence Kullback51 ; Sagawa22 ; Shiraishi23 ; Strasberg22 ; Yoshii23 and temperature; by providing analytical results. In Sec. IV, we discuss QMPE in a region with second-order EPs. Section V demonstrates the effect of complex eigenvalues on the time evolution of the density matrix leading to multiple QMPE in the observables. We summarize the results and discuss possible future directions in Sec. VI. Explicit expressions of the eigenvectors used in the analyses in Secs. IV and V, are provided in Appendix A and Appendix B respectively. We discuss QMPE briefly in region of purely exponential relaxation in Appendix C and at a third-order EP in Appendix D.
@@ II Model and protocol
$i\frac{d}{dt}|\widehat{\rho}(t)\rangle=\widehat{\mathcal{L}}|\widehat{\rho}(t)\rangle.$ (1)
@@
^$\widehat{\mathcal{L}}=\left(\begin{array}[]{cccc}1-i\tilde{\Gamma}/2&0&-\tilde{d}/2&\tilde{d}/2\\ 0&-1-i\tilde{\Gamma}/2&\tilde{d}/2&-\tilde{d}/2\\ -\tilde{d}/2&\tilde{d}/2&-i\tilde{\Gamma}&0\\ \tilde{d}/2&-\tilde{d}/2&i\tilde{\Gamma}&0\\ \end{array}\right).$ (2)
Let us clarify a few points about the matrix $\widehat{\mathcal{L}}$ at this point. Note that the dissipation $\Gamma$ makes the system non-Hermitian. More importantly, each column sum $\widehat{\mathcal{L}}$ does not equal to zero, which distinguishes Hatano’s model from the classical Markovian models studied previously Lu17 . A recent work Chatterjee_2023 showing thermal quantum Mpemba effect also deals with an effective 4-state quantum system. However, the semi-classical approach in Ref. Chatterjee_2023 considered a diagonal form of the density matrix, while the present model includes quantum coherence in the form of off-diagonal density matrix elements $\rho_{\mathrm{eg}}(t)$ and $\rho_{\mathrm{ge}}(t)$. Naturally, it is interesting to ask whether Hatano’s model, in the presence of coherence, can exhibit QMPE in observables such as temperature, energy entropy and KL divergence.
@@
The coherence terms $\rho_{\mathrm{eg}}$ and $\rho_{\mathrm{ge}}$ being complex conjugates of each other, we decompose them into real and imaginary parts as
@@ III Different regimes in $\tilde{d}-\tilde{\Gamma}$ plane and multiple QMPE
As discussed in Ref. Hatano_2019 , Hatano’s model exhibits a rich spectrum of eigenvalues. In the plane of control parameters $\tilde{d}-\tilde{\Gamma}$, it has six different regions ($\mathrm{a}_{1}$), ($\mathrm{a}_{2}$), (b), (c), (d) and (e) depending on the nature of the eigenvalues of $\widehat{\mathcal{L}}$ [Eq. (2)]. The regions ($\mathrm{a}_{1}$) and ($\mathrm{a}_{2}$) consist of complex eigenvalues, region ($\mathrm{b}$) with purely imaginary eigenvalues, regions ($\mathrm{c}$) and ($\mathrm{d}$) with second-order EPs and region ($\mathrm{e}$) containing a single third-order EP. At a $n$-th order EP, $n$ eigenvalues become equal and their eigenvectors coalesce. The regions ($\mathrm{a_{1}}$) and ($\mathrm{a_{2}}$) can be distinguished from each other depending on the largest nonzero eigenvalue which is purely imaginary in ($\mathrm{a_{1}}$) but complex in ($\mathrm{a_{2}}$). They can be also differentiated by their locations in the parameter space of $(\tilde{d},\tilde{\Gamma})$. The region ($\mathrm{a_{1}}$) is placed in relatively larger $\tilde{d}$ and smaller $\tilde{\Gamma}$, whereas the region ($\mathrm{a_{2}}$) is placed in comparatively smaller $\tilde{d}$ and larger $\tilde{\Gamma}$. All the nonzero eigenvalues in region (b) are purely imaginary and they are unequal from one another. The line (c) of second-order EPs separate region ($\mathrm{a_{1}}$) from region (b), whereas the line (d) of second-order EPs demarcate (b) from ($\mathrm{a_{2}}$). The two lines (c) and (d) of second-order EPs meet at a single point in the $\tilde{d}-\tilde{\Gamma}$ plane that defines the region (e) which is a third-order EP. See Figs. 3 and 4 in Ref. Hatano_2019 for the visual descriptions of these regions. We would like to analyze the full relaxation of the system that contains all three relaxation modes (apart from the eigenvalue zero that corresponds to the steady state) of the Lindbladian $\widehat{\mathcal{L}}$ [Eq.(2)]. The presence of multiple relaxation modes, EPs and complex eigenvalues naturally triggers the possibility of interesting features of QMPE in Hatano’s model.
@@ IV.1 Density matrix elements
where the initial density matrix elements $\rho_{n}(0)$ before quench are given by Eq. (9) and the elements $\ell_{k,n}$ are provided in Appendix A. Interestingly, in the presence of the second-order EP in region (d), the density matrix elements in Eq. (26) have picked up an additional algebraic term proportional to time $t$. This is remarkably different from the time evolution in regions (b) which has only purely exponential relaxations [see Appendix B] and region ($\mathrm{a_{1}}$) that contains damped oscillations [see Sec. V]. The linear $t$ dependence originates from the off-diagonal defect in the Jordan normal form in Eq. (22).
To investigate QMPE, we focus on the ground state density matrix element $\rho_{\mathrm{gg}}(t)$ $(\equiv\rho_{4}(t))$. To detect the QMPE, we would like to find the intersections of the two copies I and II during their relaxation, i.e. solution(s) to the equation $\Delta\rho_{\mathrm{gg}}(t)=0$. We obtain the following form of $\Delta\rho_{\mathrm{gg}}(t)$
@@ IV.5 Kullback-Leibler divergence
Figure 6: The figure (a) shows QMPE in KL divergence [Eq. (43)] where the initially higher distant copy (here I) reaches to steady state faster than the initially lower distance copy (here II). The figure (b) demonstrates the relaxation speed [Eq. (44)] of these two copies. The speed also exhibit QMPE, but the intersection time is different compared to $D-{\mathrm{KL}}(t)$. Parameters used for both figures are $\tilde{d}=4.0,\tilde{\Gamma}=\sqrt{(568+64\sqrt{2})/2},\tilde{\Gamma}_{\mathrm{I}}=20.0,\tilde{\Gamma}_{\mathrm{II}}=4.5,\tilde{d}_{\mathrm{I}}=10.0,\tilde{d}_{\mathrm{II}}=2.0$.
@@ VI Summary
Acknowledgements.- We thank Naomichi Hatano, Sosuke Ito, Ryo Hanai, Arnab Pal, Naruo Ohga and Kohei Yoshimura for useful discussions. This work is partially supported by the Grants-in-Aid for Scientific Research (Grant No. 21H01006 and No. 20K14428). A.K.C. gratefully acknowledges postdoctoral fellowship from the YITP. The numerical calculations have been done on Yukawa-21 at the YITP.
@@ D.2 Average energy
Figure 17: The figure (a) illustrates QMPE in von Neumann entropy in region (e). The difference of entropies between two copies I and II intersect zero twice, leading to double QMPE. The inset shows the second intersection that is weaker in magnitude with respect to the first intersection shown in the main figure. Parameters used in figure (a) are same as in Fig. 16. The figure (b) illustrates an example of thermal QMPE where the intersection between initially hotter and colder copies occur after the initial stage of negative temperature. The inset of figure (b) shows the full time evolution including the initial negative temperatures. Parameters used are $\tilde{d}=2\sqrt{2},\tilde{\Gamma}=6\sqrt{3},\tilde{\Gamma}_{\mathrm{I}}=14.0,\tilde{\Gamma}_{\mathrm{II}}=15.0,\tilde{d}_{\mathrm{I}}=6.0,\tilde{d}_{\mathrm{II}}=3.0$.
@@ D.4 Temperature
In this section we study the thermal QMPE in region (e). We observe thermal QMPE in region (e) by using variable dissipation protocol in Fig. 17(b). There we observe single intersection and reversal of identities (i.e. hotter becoming colder and vice versa) occur after the negative temperatures have happened. So, such temperatures can be considered as well-defined and therefore this intersection in Fig. 17(b) corresponds to single QMPE in temperature in region (e).

### v4

@@ I Introduction
Since the phenomenon of QMPE is a comparatively new area of research, it is important to explore the plausible connections of QMPE to other interesting features of quantum systems. In the present work, we would like to address some such connections. In particular, the study of exceptional points (EPs) and surfaces in non-Hermitian quantum systems and Liouvillian dynamics has caught a lot of attention in recent days Heiss12 ; Hashimoto15 ; Doppler16 ; Zhong19 ; McDonald20 ; Soleymani22 ; Ding22 ; Minganti19 ; Arkhipov20 ; Arkhipov21 ; Muller08 ; Naghiloo19 ; Saha23 ; Roccati23 ; Khandelwal21 along with their applications in quantum optics Miri19 ; Ozdemir19 . The occurrence of EPs is directly related to eigen-spectrum of a system and hence to its relaxation process. Therefore a natural curiosity would be to understand the role of EPs on the possibility of QMPE. In fact, the occurrence of double MPE (i.e. two intersections between trajectories) at an EP has already been pointed out for MPE in classical inertial suspensions Takada21a . Interestingly, the acceleration of relaxation process in open quantum systems by maximizing the gap between the zero mode and the slowest eigenmode at an EP, has recently been studied Zhou23 . In general, such speed up is desired for several application purposes and therefore it has been a subject of immense interest in both classical and quantum systems Lapolla20 ; Vu21 ; Kochsiek22 ; Ghimenti22 ; Bao22 ; Zhou23 . Since QMPE allows faster relaxation of initially higher distant trajectories, it is important to find criteria for QMPE and to understand how to utilize QMPE to achieve faster approach to steady state. Another intriguing feature in non-Hermitian quantum mechanical systems is the appearance of complex eigenvalues Bender2007 ; Moiseyev2011 ; Nakano2011 ; Hatano2015 ; Li2023 . It is noteworthy that the analysis of QMPE in the existing examples Nava19 ; Carollo21 ; Manikandan21 ; Ivander23 ; Chatterjee_2023 is based on purely exponential relaxation processes, neither the EP nor the oscillations have been considered yet. There is only an exception in Ref. Ares23 where more than one intersections between trajectories (two intersections to be precise, termed as revivals) has been observed in the entanglement entropy in the XXZ spin chain, although not discussed much and simply explained as a finite-size effect.
@@
The paper is arranged as follows. In Sec. II, we define the details of the model and the protocol used to study QMPE. The Sec. III describes different regions in the control parameter plane depending on the types of eigenvalues of the Lindbladian and the multiple QMPE is defined. In articular, we focus on regions with exceptional points and complex eigenvalues to discuss their effects on QMPE in the density matrix and observables such as energy, entropy, Kullback-Leibler (KL) divergence Kullback51 ; Sagawa22 ; Shiraishi23 ; Strasberg22 ; Yoshii23 and temperature; by providing analytical results. In Sec. IV, we discuss QMPE in a region with second-order EPs. Section V demonstrates the effect of complex eigenvalues on the time evolution of the density matrix leading to multiple QMPE in the observables. We summarize the results and discuss possible future directions in Sec. VI. Explicit expressions of the eigenvectors used in the analyses in Secs. IV and V are provided in Appendices A and B, respectively. We discuss QMPE briefly in region of purely exponential relaxation in Appendix C and at a third-order EP in Appendix D. In Appendix E, we explicitly present and discuss a diagram containing different eigenvalue regions with the variations of drive and dissipation.
@@ II Model and protocol
$\mathrm{i}\frac{d}{dt}|\widehat{\rho}(t)\rangle=\widehat{\mathcal{L}}|\widehat{\rho}(t)\rangle.$ (1)
@@
^$\widehat{\mathcal{L}}=\left(\begin{array}[]{cccc}1-\mathrm{i}\tilde{\Gamma}/2&0&-\tilde{d}/2&\tilde{d}/2\\ 0&-1-\mathrm{i}\tilde{\Gamma}/2&\tilde{d}/2&-\tilde{d}/2\\ -\tilde{d}/2&\tilde{d}/2&-\mathrm{i}\tilde{\Gamma}&0\\ \tilde{d}/2&-\tilde{d}/2&\mathrm{i}\tilde{\Gamma}&0\\ \end{array}\right).$ (2)
Let us clarify a few points about the matrix $\widehat{\mathcal{L}}$ at this point. Note that the dissipation $\Gamma$ makes the system non-Hermitian. More importantly, each column sum $\widehat{\mathcal{L}}$ does not equal to zero, which distinguishes Hatano’s model from the classical Markovian models studied previously Lu17 . A recent work Chatterjee_2023 showing thermal quantum Mpemba effect also deals with an effective 4-state quantum system. However, the semi-classical approach in Ref. Chatterjee_2023 considered a diagonal form of the density matrix, while the present model includes quantum coherence in the form of off-diagonal density matrix elements $\rho_{\mathrm{eg}}(t)$ and $\rho_{\mathrm{ge}}(t)$. Naturally, it is interesting to ask whether Hatano’s model, in the presence of coherence, can exhibit QMPE in observables such as temperature, energy, entropy and KL divergence.
@@
The coherence terms $\rho_{\mathrm{eg}}$ and $\rho_{\mathrm{ge}}$ being complex conjugates of each other, we decompose them into real (re) and imaginary (im) parts as
@@ III Different regimes in $\tilde{d}-\tilde{\Gamma}$ plane and multiple QMPE
As discussed in Ref. Hatano_2019 , Hatano’s model exhibits a rich spectrum of eigenvalues. In the plane of control parameters $\tilde{d}-\tilde{\Gamma}$, it has six different regions ($\mathrm{a}_{1}$), ($\mathrm{a}_{2}$), (b), (c), (d) and (e) depending on the nature of the eigenvalues of $\widehat{\mathcal{L}}$ [Eq. (2)]. The regions ($\mathrm{a}_{1}$) and ($\mathrm{a}_{2}$) consist of complex eigenvalues, region ($\mathrm{b}$) with purely imaginary eigenvalues, regions ($\mathrm{c}$) and ($\mathrm{d}$) with second-order EPs and region ($\mathrm{e}$) containing a single third-order EP. At a $n$-th order EP, $n$ eigenvalues become equal and their eigenvectors coalesce. The regions ($\mathrm{a_{1}}$) and ($\mathrm{a_{2}}$) can be distinguished from each other depending on the largest nonzero eigenvalue which is purely imaginary in ($\mathrm{a_{1}}$) but complex in ($\mathrm{a_{2}}$). They can be also differentiated by their locations in the parameter space of $(\tilde{d},\tilde{\Gamma})$. The region ($\mathrm{a_{1}}$) is placed in relatively larger $\tilde{d}$ and smaller $\tilde{\Gamma}$, whereas the region ($\mathrm{a_{2}}$) is placed in comparatively smaller $\tilde{d}$ and larger $\tilde{\Gamma}$. All the nonzero eigenvalues in region (b) are purely imaginary and they are unequal from one another. The line (c) of second-order EPs separate region ($\mathrm{a_{1}}$) from region (b), whereas the line (d) of second-order EPs demarcate (b) from ($\mathrm{a_{2}}$). The two lines (c) and (d) of second-order EPs meet at a single point in the $\tilde{d}-\tilde{\Gamma}$ plane that defines the region (e) which is a third-order EP. See Appendix E for the visual descriptions of these regions (also see Figs. 3 and 4 in Hatano_2019 ). We would like to analyze the full relaxation of the system that contains all three relaxation modes (apart from the eigenvalue zero that corresponds to the steady state) of the Lindbladian $\widehat{\mathcal{L}}$ [Eq.(2)]. The presence of multiple relaxation modes, EPs and complex eigenvalues naturally triggers the possibility of interesting features of QMPE in Hatano’s model.
@@ IV.1 Density matrix elements
where the initial density matrix elements $\rho_{n}(0)$ before quench are given by Eq. (9) and the elements $\ell_{k,n}$ are provided in Appendix A. Interestingly, in the presence of the second-order EP in region (d), the density matrix elements in Eq. (26) have picked up an additional algebraic term proportional to time $t$. This is remarkably different from the time evolution in regions (b) which has only purely exponential relaxations [see Appendix C] and region ($\mathrm{a_{1}}$) that contains damped oscillations [see Sec. V]. The linear $t$ dependence originates from the off-diagonal defect in the Jordan normal form in Eq. (22).
To investigate QMPE, we focus on the ground state density matrix element $\rho_{\mathrm{gg}}(t)$ $(\equiv\rho_{4}(t))$. To detect the QMPE, we would like to find the intersections of the two copies I and II during their relaxation, i.e. solution(s) to the equation $\Delta\rho_{\mathrm{gg}}(t)=\rho_{\mathrm{gg}}^{\mathrm{I}}(t)-\rho_{\mathrm{gg}}^{\mathrm{II}}(t)=0$, where $\rho_{\mathrm{gg}}^{\mathrm{I}}(t)$ and $\rho_{\mathrm{gg}}^{\mathrm{II}}(t)$ are ground state probabilities for copies I and II, respectively. We obtain the following form of $\Delta\rho_{\mathrm{gg}}(t)$
@@ IV.5 Kullback-Leibler divergence
Figure 6: The figure (a) shows QMPE in KL divergence [Eq. (43)] where the initially higher distant copy (here I) reaches to steady state faster than the initially lower distance copy (here II). The figure (b) demonstrates the relaxation speed [Eq. (44)] of these two copies. The speed also exhibit QMPE, but the intersection time is different compared to $D_{\mathrm{KL}}(t)$. Parameters used for both figures are $\tilde{d}=4.0,\tilde{\Gamma}=\sqrt{(568+64\sqrt{2})/2},\tilde{\Gamma}_{\mathrm{I}}=20.0,\tilde{\Gamma}_{\mathrm{II}}=4.5,\tilde{d}_{\mathrm{I}}=10.0,\tilde{d}_{\mathrm{II}}=2.0$.
@@ VI Summary
Acknowledgements.- We thank Naomichi Hatano, Sosuke Ito, Ryo Hanai, Arnab Pal, Naruo Ohga and Kohei Yoshimura for useful discussions. This work is partially supported by the JSPS KAKENHI (Grants No. 21H01006 and No. 20K14428). A.K.C. gratefully acknowledges a postdoctoral fellowship from the YITP. For numerical calculations, we have used the computing facilities comp .
@@ D.2 Average energy
Figure 17: The figure (a) illustrates QMPE in von Neumann entropy in region (e). The difference of entropies between two copies I and II intersect zero twice, leading to double QMPE. The inset shows the second intersection that is weaker in magnitude with respect to the first intersection shown in the main figure. Parameters used in figure (a) are same as in Fig. 16. The figure (b) illustrates an example where the intersection between initially hotter and colder copies occurs before the initial stage of negative temperature. The inset of figure (b) shows the full time evolution including the initial negative temperatures. Parameters used are $\tilde{d}=2\sqrt{2},\tilde{\Gamma}=6\sqrt{3},\tilde{\Gamma}_{\mathrm{I}}=14.0,\tilde{\Gamma}_{\mathrm{II}}=15.0,\tilde{d}_{\mathrm{I}}=6.0,\tilde{d}_{\mathrm{II}}=3.0$.
@@ D.4 Temperature
In this section we study the thermal QMPE in region (e). We investigate thermal QMPE in region (e) by using variable dissipation protocol in Fig. 17(b). There we observe single intersection and reversal of identities (i.e. hotter becoming colder and vice versa) occur before the negative temperatures have happened. Note that this is contrary to the thermal QMPEs obtained in Figs. 5, 7 and 14(a) where we obtain single or multiple intersections after the initial stage of negative temperatures have passed. So, such temperatures may not be considered as well-defined and therefore this intersection in Fig. 17(b) may not correspond to single QMPE in temperature in region (e). We have to further look for other parameters’ values which can lead us to intersections after the initial stage of negative temperatures so that thermal QMPE can be obtained unambiguously like Figs. 5, 7 and 14(a).